A tank can hold eighteen litres of water. Tap A takes six minutes to fill the tank, Tap B takes three minutes to fill the tank . If the two taps are turned on at the same time, how long would they take to fill the tank?
Answer by nerdybill(7384)

A tank can hold eighteen litres of water. Tap A takes six minutes to fill the tank, Tap B takes three minutes to fill the tank . If the two taps are turned on at the same time, how long would they take to fill the tank?
.
Let x = time (minutes) it takes for both
then
x(1/6 + 1/3) = 1
multiplying both sides by 18:
x(3 + 6) = 18
x(9) = 18
x = 18/9
x = 2 minutes